If my theoretical yield is 75 g of AI2O3 and I actually get 20 g of AI2O3 in the lab, what is my percent yield?
1
Expert's answer
2020-02-25T04:53:50-0500

ω=(m1/m2) *100%

ω - yield

m1 - obtained (actual mass obtained in the lab)

m2 - theoretical mass

ω=(20/75) * 100% = 26.67%